### v2

# Implications of portal vector-like lepton on associated Higgs production at a multi-TeV muon collider

Abstract
We have explored a portal vector-like lepton (pVLL) extension of the Standard Model (SM) and studied its implications for Higgs and vector-boson associated production ($hV$, with $V=Z$-boson or dark photon) at a future muon collider facility. We show that while the $\mu^{+}\mu^{-}\to hZ$ production rate remains close to its SM prediction in a wide range of parameter space, the rate for $\mu^{+}\mu^{-}\to h\gamma_{d}$ can be substantially enhanced owing to the non-decoupling nature of the interaction involving the heavy lepton, the muon and the dark photon. We demonstrate that the $h\gamma_{d}$ production rate can exceed the corresponding $hZ$ rate by a factor of 11100100 within the perturbative unitarity limit, making it a promising channel for probing Higgs interactions and potential new physics effects.
Furthermore, this process can also be used to constrain the dark photon mass ($m_{\gamma_{d}}$) and/or the dark gauge coupling ($g_{d}$) consistent with the current muon g2g-2 measurements within the pVLL framework. We perform a detailed collider analysis of the $h\gamma_{d}$ process in the $b\bar{b}\penalty 10000\ +$ missing energy final state. A $2\sigma$ exclusion limit for $m_{\gamma_{d}}$ up to $80$ GeV is obtained assuming $g_{d}=0.05$, $\sin\theta_{L}=4\times 10^{-5}$, $\sin\theta_{s}=0.05$, for a heavy lepton mass $\penalty 10000\ \sim 3$ TeV at a $3$ TeV muon collider with an integrated luminosity of $1$ ab-1.
1 Introduction
The discovery of the Higgs boson [1, 2] marks a remarkable triumph of theoretical and experimental particle physics. To probe its properties with greater precision, however, we require an abundant production of Higgs bosons at high-energy (HE) and high-luminosity (HL) facilities such as the Large Hadron Collider (LHC) and future collider experiments. Higgs production in association with a gauge boson is a well-studied process at lepton colliders [3, 4, 5, 6, 7], both in past experiments [8, 9] and in proposed future facilities [10, 11, 12, 13]. The presence of new physics, particularly, scenarios predicting vector-like leptons (VLLs), can significantly alter the associated Higgs production cross section at lepton colliders—either through modified Higgs couplings, the appearance of new diagrams, or both. Moreover, such scenarios may open up additional Higgs production channels with enhanced rates. In this work, we explore the implications of VLLs acting as portal matter on the Higgs production cross section at a future muon collider.
VLLs constitute a simple yet well-motivated extension of the Standard Model (SM) of particle physics. VLL extension of the SM has been studied extensively in the literature in the context of various anomalies, including the muon g2g-2, the $W$-boson mass, the Cabibbo angle anomaly, and dark matter phenomenology [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. They naturally arise in a variety of theoretical frameworks [28, 29, 30, 31, 32, 33], including grand unified theories, models addressing the Higgs mass hierarchy problem, and constructions inspired by string theory. The search for VLLs at collider experiments remains an active area of investigation, including both prompt and long-lived signatures, primarily in SM final states [34, 35, 36]. VLL is also a potential candidate to search for at the future colliders such as future $pp,\penalty 10000\ e^{+}e^{-}$, muon collider [37, 38, 39, 40, 41].
In recent years, the idea of VLLs as portal matter [42, 43, 44, 18] or portal VLLs (pVLLs) has gained significant interest owing to their rich phenomenology, their ability to evade existing collider constraints, and their potential connection to the dark sector, particularly dark matter phenomenology [45, 46].
In this work, we consider an extension of the Standard Model (SM) with an $SU(2)$-singlet VLL that transforms under both the SM hyper-charge group $U(1)_{Y}$ and an additional abelian dark gauge group $U(1)_{D}$ [42, 43, 44, 18]. The gauge boson associated with $U(1)_{D}$ is referred to as the dark photon ($\gamma_{d}$). The model also contains an additional complex scalar field, singlet under the SM gauge group but carrying non-zero $U(1)_{D}$ charge. The full gauge symmetry, $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}\times U(1)_{D}$, is spontaneously broken to $SU(3)_{C}\times U(1)_{\rm EM}$ through the non-vanishing vacuum expectation values (VEVs) of the SM Higgs doublet and the dark scalar. The VEV of the dark scalar not only generates the dark photon mass but also induces mixing between the pVLL and SM leptons through a Yukawa-type portal interaction. The $U(1)_{Y}$ hypercharge assignment of the pVLL considered here corresponds to an electric charge of 1-1, and it is allowed to mix only with one generation of SM charged leptons in order to evade stringent bounds from lepton-flavor violation. In this work, we focus on pVLL mixing with the second-generation charged lepton.
An important feature of this model is that the right-handed second generation charged lepton ($\mu_{R}^{\prime}$) and the pVLL ($\mu_{p}^{\prime}$) carry identical $SU(2)_{L}\times U(1)_{Y}$ charges. Consequently, any rotation between $\mu_{R}^{\prime}$ and $\mu_{p_{{}_{R}}}^{\prime}$ preserves the diagonal structure of the $Z$-boson couplings, ensuring that no $\mu_{R}-\mu_{p_{{}_{R}}}-Z$ interaction arises after transforming to the mass basis. In contrast, since the muon is uncharged under $U(1)_{D}$, such a rotation induces an off-diagonal $\mu_{R}-\mu_{p_{{}_{R}}}-\gamma_{d}$ coupling, suppressed by $\sin\theta_{R}$, the sine of the corresponding mixing angle. One can have a sizable value of $\sin\theta_{R}\sim 0.1$ for heavy muon mass $\sim$ a few TeV even when the mixing angle ($\sin\theta_{L}$) involving the corresponding left-chiral fermions is kept at a tiny value $\sim 10^{-5}$.
We explore this non-decoupling effect within the limit of perturbative unitarity at the proposed future muon collider, focusing on Higgs production in association with a vector boson ($\mu^{+}\mu^{-}\to hV$ with $V=Z,\gamma_{d}$). The muon collider [47, 48, 49] is envisioned as a high-energy circular lepton collider that combines the clean experimental environment of lepton collisions with access to multi-TeV center-of-mass energies ($\sqrt{s}$), in contrast to circular $e^{+}e^{-}$ machines. It is designed to operate at energy stages from an initial 3 TeV to 10 TeV (or higher), corresponding to circumferences of about 4.5 km and 10 km, respectively, with expected peak integrated luminosities of 1 ab-1 at $\sqrt{s}=3$ TeV and 10 ab-1 at $\sqrt{s}=10$ TeV [50].
We have shown that while the $hZ$ production cross section remains essentially unchanged from the SM prediction, the $h\gamma_{d}$ production rate can be significant even for very small $\sin\theta_{L}$, owing to the non-decoupling effect. In this scenario, the dark photon is treated as an invisible particle [^1] This will be the case if the dark photon couples to dark matter particle and predominantly decays to invisible final state.. We analyze this process in the $b\bar{b}$ plus missing energy final state at $\sqrt{s}=3$ and $10$ TeV muon colliders, where the $b\bar{b}$ system is highly boosted due to the $2\to 2$ kinematics and large center-of-mass energy. The dominant SM backgrounds to this channel are processes leading to $b\bar{b}\nu\bar{\nu}$ final state, with $h\nu\bar{\nu}$ and $Z\nu\bar{\nu}$ accounting for more than $90\%$ of the total cross section. We demonstrate that the requirement of jet substructure, together with additional kinematic selections such as missing energy and the $b\bar{b}$ invariant mass, can efficiently suppress these backgrounds.
The main highlight of this work is that the $h\gamma_{d}$ process not only yields a higher Higgs production rate than $hZ$ production at a muon collider, but also offers a noble way to explore dark photons in the mass range 1010100100 GeV, independent of the gauge kinetic mixing parameter ($\varepsilon$) [51, 52, 53, 54]. This is because of the fact that the muon–heavy muon–Higgs vertex is sensitive to the dark photon mass ($m_{\gamma_{d}}$) and the dark gauge coupling ($g_{d}$). Consequently, the absence of any observable enhancement in Higgs production along with missing energy at a muon collider would imply a stringent limit on $m_{\gamma_{d}}$ or $g_{d}$, for fixed values of other relevant model parameters. We illustrate this using several representative benchmark points and show that, even for fermion mixing angle as small as $\sim 10^{-6}$ and heavy muon masses in the range 1133 TeV, a wide range of dark photon masses can be excluded. We illustrarte that a substantial region of the parameter space consistent with current measurement of muon g2g-2 can be probed at the future muon collider using our analysis.
The remainder of the manuscript is organized as follows. In Section 2, we present the theoretical framework considered in this work to illustrate the impact of portal VLLs on associated Higgs production at muon collider, along with a discussion of relevant constraints on the model parameters. Section 3 describes the collider analysis of Higgs production in association with a vector boson. The results of our analysis are presented in Section 4, and finally we conclude in Section 5.
2 Theoretical Framework
We consider a framework in which the Standard Mode is extended by a vector-like lepton, singlet under $SU(3)_{C}\times SU(2)_{L}$ with $U(1)_{Y}$ hypercharge -$1$. The VLL plays the role of a portal matter, linking the visible SM sector with a dark sector. A simplified realization of such a scenario is to have a VLL transforming non-trivially under an additional local $U(1)_{D}$ symmetry of the dark sector with corresponding gauge boson referred to as the dark photon ($\gamma_{d}$). The dark photon acquires mass through spontaneous breaking of $U(1)_{D}$, achieved by introducing a complex scalar field $\Phi_{d}$ which is a singlet under the SM gauge group but charged under $U(1)_{D}$. Furthermore, if $\Phi_{d}$ carries the same $U(1)_{D}$ charge as the pVLL, a Yukawa-type portal interaction $\omega_{f}\,\Phi_{d}\,\overline{\mu}^{\prime}_{p_{L}}\mu^{\prime}_{R}$ is permitted. The Yukawa interaction reduces to a bilinear term between the pVLL and the SM lepton once $\Phi_{d}$ gets a non-zero VEV. In this work, we focus on the case where the pVLL mixes with the second-generation charged lepton only to avoid constraints coming from flavor violating observables. The assignment of charge of the relevant field content is presented in Table 1.
Fields $SU(3)_{C}$ $SU(2)_{L}$ $Y$ $Y_{d}$
$\mu^{{}^{\prime}}_{R}$ 1 1 -1 0
$L_{L}=\begin{pmatrix}\nu_{\mu_{L}}\\ \mu^{{}^{\prime}}_{L}\end{pmatrix}$ 1 2 -1/2 0
$\Phi$ 1 2 1/2 0
$\mu^{\prime}_{p}$ 1 1 -1 1
$\Phi_{d}$ 1 1 0 1
Table 1: Representation of relevant fields and their $U(1)$ charge assignment under the extended gauge group.
In Table 1, $\mu_{R}^{\prime}$ and $L_{L}$ represent the second-generation SM right-handed charged lepton and left-handed lepton doublet, respectively, and $\Phi$ corresponds to the SM complex scalar field doublet. The additional VLL singlet under $SU(2)_{L}$ is represented by $\mu_{p}^{\prime}$. $\Phi_{d}$ is the complex scalar field that does not carry charges under the SM gauge group. All fields are expressed in their interaction basis. The SM Lagrangian now will be extended with new terms coming from the effect of the BSM interactions we have considered in our model.
The complete Lagrangian of the model invariant under $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}\times U(1)_{D}$ can be expressed as
$\mathcal{L}$ $=$ $\mathcal{L}_{\rm Gauge}+\mathcal{L}_{\rm Scalar}+\mathcal{L}_{\rm Fermions}+\mathcal{L}_{\rm Yukawa}$ (1)
where $\mathcal{L}_{\rm Gauge},\penalty 10000\ \mathcal{L}_{\rm Scalar},\penalty 10000\ \mathcal{L}_{\rm Fermions}$ are part of the full Lagrangian containing gauge invariant kinetic terms involving the gauge fields, scalar fields including corresponding scalar potential, and fermion fields including the mass term of the pVLL, respectively. Finally, $\mathcal{L}_{\rm Yukawa}$ represents the trilinear Yukawa interactions involving scalar and fermions.
In particular, the gauge sector of the theory is given by
$\mathcal{L}_{\rm Gauge}$ $=$ $-\frac{1}{4}G_{\mu\nu}^{a}G^{a,\mu\nu}-\frac{1}{4}W_{\mu\nu}^{i}W^{i,\mu\nu}-\frac{1}{4}B_{\mu\nu}B^{\mu\nu}-\frac{1}{4}B_{d,\mu\nu}B_{d}^{\mu\nu}+\frac{\varepsilon}{2\cos{\theta_{w}}}B_{d,\mu\nu}B^{\mu\nu},$ (2)
where $G_{\mu\nu}^{a}$ denotes the QCD field strength tensor with $a=1,\dots,8$ labeling the adjoint of $SU(3)_{C}$, $W_{\mu\nu}^{i}$ ($i=1,\penalty 10000\ 2,\penalty 10000\ 3$) are the field strength tensors of $SU(2)_{L}$, $B_{\mu\nu}$ corresponds to the $U(1)_{Y}$ hypercharge, and $B_{d,\mu\nu}$ is the field strength tensor of the dark $U(1)_{D}$.
The scalar-sector Lagrangian is
$\mathcal{L}_{\rm Scalar}$ $=$ $|D_{\mu}\Phi|^{2}+|D_{\mu}\Phi_{d}|^{2}-V(\Phi,\Phi_{d})\,$, (3)
with the scalar potential
$V(\Phi,\Phi_{d})$ $=$ $-\mu_{h}^{2}|\Phi|^{2}-\mu_{h_{d}}^{2}|\Phi_{d}|^{2}+\lambda_{h}|\Phi|^{4}+\lambda_{h_{d}}|\Phi_{d}|^{4}+\lambda_{hh_{d}}|\Phi|^{2}|\Phi_{d}|^{2}\,.$ (4)
The symmetry breaking condition and boundedness of the potential require $\mu_{h}^{2},\mu_{h_{d}}^{2}>0$ and $(\lambda_{h}\lambda_{h_{d}}-\lambda_{hh_{d}}^{2}),\lambda_{h},\lambda_{h_{d}}>0$. The field configurations of $\Phi$ and $\Phi_{d}$ at the minima of the scalar potential in unitary gauge is given by
$\Phi=\begin{pmatrix}0\\ \dfrac{v_{{}_{\rm EW}}+h_{1}}{\sqrt{2}}\end{pmatrix},\quad\Phi_{d}=\dfrac{1}{\sqrt{2}}(v_{d}+h_{2}).$ (5)
where $h_{1}$ and $h_{2}$ denote the fluctuation around the minimum field configurations and $v_{{}_{\rm EW}}$ and $v_{d}$ represent the VEV of the respective scalar fields [^2] The mass of the dark photon accordingly is given by $m_{\gamma_{d}}=g_{d}v_{d}$.. We denote the fields in the mass basis by $h$ and $h_{d}$ which correspond to the observed Higgs and the hypothetical dark Higgs fields, respectively. The physical scalar fields ($h,\penalty 10000\ h_{d}$) are related to the ($h_{1},\penalty 10000\ h_{2}$) by the following unitary transformation
$\begin{pmatrix}h\\ h_{d}\end{pmatrix}$ $=$ $\begin{pmatrix}\cos\theta_{s}&-\sin\theta_{s}\\ \sin\theta_{s}&\cos\theta_{s}\end{pmatrix}\begin{pmatrix}h_{1}\\ h_{2}\end{pmatrix}$ (6)
Here, $\theta_{s}$ is the scalar mixing angle.
The expression for $\mathcal{L}_{\rm Fermion}$ and $\mathcal{L}_{\text{Yukawa}}$ appearing in Eq. (1) involving second generation leptons and pVLL consistent with gauge symmetry are given by
[…]
$\mathcal{L}_{\text{Yukawa}}$ $\supset$ $-{y_{m}\,\bar{L}_{\mu}\Phi\,\mu_{R}^{\prime}}-\omega_{f}\,\Phi_{d}\,\overline{\mu}^{\prime}_{p_{L}}\mu^{\prime}_{R}+\mathrm{H.c.}$ (8)
The mixing between $\mu^{\prime}$ and $\mu^{\prime}_{p}$ is generated once the dark Higgs assumes non-zero VEV. After symmetry breaking, the mass matrix in the interaction basis $\mu^{\prime}$ and $\mu^{\prime}_{p}$ takes the form
$\mathcal{M}=\begin{pmatrix}\dfrac{y_{m}\,v_{{}_{\rm EW}}}{\sqrt{2}}&0\\[8.5359pt] \dfrac{\omega_{f}\,v_{d}}{\sqrt{2}}&m_{\mu^{\prime}_{p}}\end{pmatrix}$ (9)
The mass eigenbasis $\mu$ and $\mu_{p}$ in which the mass matrix is diagonal, represent the physical muon and the corresponding heavy muon partner, respectively. The mass eigenbasis and the interaction basis are connected by the following bi-unitary transformation,
$\begin{pmatrix}\mu\\ \mu_{p}\end{pmatrix}_{{L,R}}=\begin{pmatrix}\cos{\theta_{{}_{L,R}}}&-\sin{\theta_{{}_{L,R}}}\\ \sin{\theta_{{}_{L,R}}}&\cos{\theta_{{}_{L,R}}}\end{pmatrix}\begin{pmatrix}\mu^{\prime}\\ \mu_{p}^{\prime}\end{pmatrix}_{{L,R}}$ (10)
where, $\theta_{{}_{L,R}}$} is the mixing angle between the left (right) chiral muon and the corresponding muon partner with same chirality.
The model parameters $y_{m},\penalty 10000\ \omega_{f},\penalty 10000\ m_{\mu^{\prime}_{p}}$ appearing in the Lagrangian of Eq. (7) and (8) can be expressed in terms of the physical parameters $m_{\mu},\penalty 10000\ m_{\mu_{{}_{P}}}\penalty 10000\ \text{and}\penalty 10000\ \sin\theta_{L}$ as
$y_{m}$ $=$ $\dfrac{\sqrt{2}}{v_{{}_{\text{EW}}}}m_{{}_{\mu_{p}}}D$ (11a)
$\omega_{f}$ $=$ $\dfrac{m_{{}_{\mu_{p}}}(1-\dfrac{m_{\mu}^{2}}{m_{{}_{\mu_{p}}}^{2}})}{\sqrt{2}v_{d}D}\sin 2\theta_{L}$ (11b)
$m_{{}_{\mu^{\prime}_{p}}}$ $=$ $\dfrac{m_{\mu}}{D}$ (11c)
where $D$ is a dimensionless quantity given by $D=\sqrt{\dfrac{m_{{\mu}}^{2}}{m_{{}_{\mu_{p}}}^{2}}\cos^{2}{\theta_{L}}+\sin^{2}{\theta_{L}}}$.
The absolute value of the sine of the mixing angle between the left chiral second generation charged lepton and the pVLL is expressed as
$\left|\sin\theta_{L}\right|=\frac{1}{2}\sqrt{\frac{2m_{{}_{\mu_{p}}}^{2}-2m_{\mu}^{2}-\omega_{f}^{2}v_{d}^{2}}{m_{{}_{\mu_{p}}}^{2}-m_{\mu}^{2}}\left(1-\sqrt{1-\frac{8\omega_{f}^{2}v_{d}^{2}m_{\mu}^{2}}{\left(2m_{{}_{\mu_{p}}}^{2}-2m_{\mu}^{2}-\omega_{f}^{2}v_{d}^{2}\right)^{2}}}\right)}$ (12)
The reality of $\sin\theta_{L}$ requires
$\left|\omega_{f}\right|<\sqrt{2}\penalty 10000\ \dfrac{(m_{{}_{\mu_{p}}}-m_{\mu})}{v_{d}}$ (13)
On the other hand, the perturbative unitarity [55] demands
$\left|\omega_{f}\right|<4\penalty 10000\ \sqrt{2\pi}$ (14)
These two conditions can be integrated as
$\left|\omega_{f}\right|<\sqrt{2}\penalty 10000\ \min\left(\frac{m_{{}_{\mu_{p}}}-m_{\mu}}{v_{d}},4\sqrt{\pi}\right)$ (15)
This framework naturally gives rise to three portal interactions connecting the visible and dark sectors: the gauge kinetic mixing, Yukawa, and scalar portals. In particular, gauge kinetic mixing between the SM $U(1)_{Y}$ and dark $U(1)_{D}$ arises at one loop and is strongly constrained by low-energy precision measurements [53, 56, 57], with allowed values sensitive to the dark photon mass. The dark photon couples to the SM fermions exclusively via gauge kinetic mixing, except for muons, which also couple through the muon–heavy muon mixing. While the kinetic-mixing portal doesn’t play any crucial role, the scalar and Yukawa portals significantly affect Higgs-associated production at a muon collider in this scenario.
The list of free parameters relevant for the discussions in the later sections are mass of the heavy muon ($m_{\mu_{{}_{p}}}$), mixing angle between the muon and its heavier counter part ($\sin\theta_{L}$), mass of the dark photon ($m_{\gamma_{d}}$), the dark gauge coupling ($g_{d}$), and the scalar mixing angle ($\sin\theta_{s}$) along with the SM parameters. Mass of the additional scalar ($m_{h_{d}}$) is an independent parameter of the model, however, its impact on our analysis is less significant unlike the scalar mixing angle ($\sin\theta_{s}$). The kinetic mixing parameter $\varepsilon$ also does not enter directly in our analysis except the requirement that its value should be small enough to evade the existing constraints and the dark photon decays exclusively to invisible mode via $U(1)_{D}$ gauge coupling in the range of $m_{\gamma_{d}}$ relevant for our analysis.
2.1 Constraints
In this subsection we discuss various observational constraints on the allowed values of the relevant model parameters discussed above.
2.1.1 Constraint on the scalar mixing angle
The scalar mixing angle which simply plays a role of overall coupling rescaling factor in the analysis presented in this paper is constrained by various observations across the mass of the heavy scalar [58, 59, 60, 61, 62, 63, 64]. For scalar mass below $100$ GeV there are constraints from the LEP experiment and invisible decay of the Higgs boson. For example, if $\penalty 10000\ m_{h_{d}}<\frac{m_{h}}{2}\penalty 10000\$the corresponding bound on the scalar mixing angle coming from the measurement of the Higgs to invisible final state at LHC is of the order of $0.01$ for $v_{d}\sim 100$ GeV[65, 66]. The LEP experiment suggests $\sin{\theta_{s}}$ in the range (0.90.9980.9-0.998) is ruled out for $m_{h_{d}}$ in the range 1010010-100 GeV and $v_{d}\sim v_{{}_{\rm EW}}$ [58, 59, 60, 61, 63]. For masses beyond $100$ GeV the constraint coming from collider experiment LHC experiment is more relevant. The constraint on $\sin\theta_{s}$ in the range $\penalty 10000\ 100\penalty 10000\ \text{GeV}<m_{h_{d}}<200\penalty 10000\ \text{GeV}\penalty 10000\$is primarily due to the Higgs signal strength measurement and $\lvert\sin\theta_{s}\rvert>0.1$ is ruled out in this mass range [63] with $v_{d}\sim 0.1\times v_{{}_{\rm EW}}$. For $\penalty 10000\ m_{h_{d}}>200\penalty 10000\$GeV, $\penalty 10000\ \sin{\theta_{s}}>0.2\penalty 10000\$is disfavored by the LHC data [61, 67].
2.1.2 Lepton flavor universality constraint
In the SM, the universality of the gauge coupling ensures that the $Z$ boson decays in the leptonic final states, in particular, in the final state with electron or muon with equal rate and the ratio of these two rates is predicted to be close to unity in the SM [19]. This has been verified in the LEP experiment at CERN and the observed value of the ratio of these two decay rates ($R$) is found to be [68]
$R$ $=$ ${\Gamma(Z\to\mu^{+}\mu^{-})}/{\Gamma(Z\to e^{+}e^{-})}=1.0009\pm 0.0028$ (16)
The model considered in this work introduces a mixing between the second-generation charged lepton and the pVLL. As a consequence, the $Z-\mu-\mu$ vertex, in particular, the $Z-\mu^{-}_{{}_{L}}-\mu^{-}_{{}_{L}}$ interaction vertex is modified to $\dfrac{g_{z}}{2}(2\sin^{2}\theta_{W}-\cos^{2}\theta_{L})$ (See Eq. (A.7) in Appendix) compared to that [^3] The corresponding factor for the SM $Z-\mu^{-}_{{}_{L}}-\mu^{-}_{{}_{L}}$ vertex is $\dfrac{g_{z}}{2}(2\sin^{2}\theta_{W}-1)$ in the SM while corresponding right-handed vertex factor remains the same due to the choice of our charge assignment. Therefore, for large mixing one expects deviation in the predicted value of $R$ inconsistent with the above observation. It turns out that, the lepton flavor universality constraint strongly disfavor ($\sin{\theta_{L}}>0.030(0.047)$) at $1\sigma$ ($2\sigma$) level [27].
2.2 LHC constraints
Various extensions of the SM containing the VLLs have been probed at the LHC as well as at the LEP experiment. Due to the limited center-of-mass energy available at the LEP collider, VLLs with mass above $102.6\penalty 10000\ \text{GeV}$ [69] are not excluded by the LEP experiment. VLL decaying in the SM mode i.e, $\nu W,\penalty 10000\ \ell Z$ has been probed at the LHC in the multi-lepton [34, 70] and hadronically decaying tau lepton with additional jets and/or missing transverse energy channels [71, 72]. The ATLAS experiment has searched for VLL mixing with first two generations of charged leptons and has obtained an exclusion limit up to 320 GeV (1220 GeV) on the allowed values of the singlet (doublet) -like electron leptons [34]. The corresponding bound for VLL mixing with muons is 400 GeV (1270 GeV) obtained by same experiment if it is an $SU(2)_{L}$ singlet (doublet). The CMS experiment at the LHC sets a lower limit on the allowed value of the doublet (singlet) vector-like tau lepton mass to be $\gtrsim 1045\penalty 10000\ \text{GeV}\penalty 10000\ (150\penalty 10000\ \text{GeV})$ with 95 % Confidence Level [70] using $138\penalty 10000\ \mathrm{fb}^{-1}$ data collected during 2016–2018. The corresponding lower limit obtained by ATLAS experiment on the allowed value of mass of the tau VLL transforming as $SU(2)_{L}$ doublet is $900$ GeV [71]. Moreover, these limits can be relaxed further if the VLLs are allowed to decay in the non-standard modes such as pVLL decaying to a lepton and a dark photon as considered in the present model.
2.3 Electroweak precision constraints
The presence of VLL has various observable consequences and can potentially modify the predictions of the SM. It can contribute to the modification of muon life-time, the branching ratio of $Z\to\mu^{+}\mu^{-}$, total width of the Z and W bosons and correction to the oblique parameters etc. Since the VLL carries electromagnetic charge it can modify the Higgs to di-photon decay rate as well. In our model, the pVLL can couple to the Higgs via muon-heavy muon mixing and/or Higgs-dark Higgs mixing. Therefore, the corresponding interaction vertices will be proportional to $\sim(y_{m}\cos\theta_{R}\cos\theta_{L}\cos\theta_{s}\penalty 10000\ +\penalty 10000\ \omega_{f}\cos\theta_{R}\sin\theta_{L}\allowbreak\sin\theta_{s}$). Hence, for small $\sin\theta_{L}$ and $\sin\theta_{s}$ the deviation from the SM $h-\mu-\mu$ vertex factor is less significant. The correction to the oblique parameters, in particular, S and T due to the singlet VLL is expressed as [73]
$\Delta T$ $=$ $\frac{m_{\mu}^{2}}{16\pi\sin^{2}_{W}m_{W}^{2}}\sin^{2}\theta_{L}\left[-(1+\cos^{2}\theta_{L})+2\cos^{2}\theta_{L}\frac{r_{\mu}}{r_{\mu}-1}\log(r_{\mu})+r_{\mu}\sin^{2}\theta_{L}\right]$ (17)
[…]
$\left.\left.\hskip 56.9055pt+\frac{3(r_{\mu}+1)\left(r^{2}_{\mu}-4r_{\mu}+1\right)}{\left(1-r_{\mu}\right)^{3}}\log(r_{\mu})\right)\right]$
where, $r_{\mu}=m_{{}_{\mu_{p}}}/m_{\mu}$.
The updated global fit to the the electroweak precision observables tells that these corrections are constrained as follows [74]:
$S$ $=$ $-0.05\pm 0.07$ (19)
$T$ $=$ $0.00\pm 0.06$ (20)
In the remaining of this paper, we work with benchmark points consistent with all the constraints mentioned above.
2.4 Muon g2g-2
The presence of muon–heavy muon–dark photon vertex can give leading contribution to the anomalous muon g2g-2 at one-loop in this pVLL framework. In the limit $m_{\mu_{p}}\approx m_{\mu^{\prime}_{p}}$ and $m_{\mu}\approx m_{\mu^{\prime}}=\dfrac{1}{\sqrt{2}}y_{m}v_{{}_{\rm EW}}$, this contribution is given by [21, 27]
$\Delta a_{\mu}$ $\approx$ $\begin{cases}\penalty 10000\ \penalty 10000\ \dfrac{g^{2}_{d}\,m_{\mu^{\prime}_{p}}\,m_{\mu}}{16\pi^{2}\,m^{2}_{\gamma_{d}}}\sin 2\theta_{R}\,\sin 2\theta_{L},&\penalty 10000\ m_{\mu^{\prime}_{p}}>>m_{\gamma_{d}}\\ \penalty 10000\ \penalty 10000\ \dfrac{g^{2}_{d}\,m_{\mu^{\prime}_{p}}\,m_{\mu}}{4\pi^{2}\,m^{2}_{\gamma_{d}}}\sin 2\theta_{R}\,\sin 2\theta_{L},&\penalty 10000\ m_{\mu}<<m_{\mu^{\prime}_{p}}<<m_{\gamma_{d}}\end{cases}$ (21)
The current world average value of $\Delta a_{\mu}=a^{\rm exp}_{\mu}-a^{\rm SM}_{\mu}$, combining Fermilab Run 1–6 data [75] with the results of Brookhaven experiment [76] is
[…]
which corresponds to $0.6\sigma$ deviation compared to the SM prediction ($a^{\rm SM}_{\mu}=0.00116592033(62)$ [77]).
In Fig. 1, we show the muon g2g-2 allowed region of parameter space both in the $m_{\gamma_{d}}-g_{d}$ and $m_{\mu_{p}}-m_{\gamma_{d}}$ planes consistent with $\Delta a_{\mu}$ mentioned in Eq. (21) at $1\sigma$ level. Fig LABEL:subfig:muon_gm2_mzp_vs_gd corresponds to $\sin\theta_{L}=4\times 10^{-5}$ and four different heavy muon masses (1 TeV, 1.5 TeV, 2 TeV and 3 TeV), whereas Fig LABEL:subfig:muon_gm2_constraint_mzp_vs_mmup corresponds to two distinct sets of $\left(g_{d},\penalty 10000\ \sin\theta_{L}\right)$ with values $\left(0.05,\penalty 10000\ 4\times 10^{-5}\right)$ and $\left(0.5,\penalty 10000\ 1.66\times 10^{-6}\right)$.
Figure 1: Region (coloured) of parameter space allowed by muon g2g-2 measurement at $1\sigma$ level [77] in the (a) $m_{\gamma_{d}}-g_{d}$ plane for $\sin\theta_{L}=4\times 10^{-5}$ and four different $m_{\mu_{p}}$ values (11 TeV, $1.5$ TeV, 22 TeV and 33 TeV) and (b) $m_{\mu_{p}}-m_{\gamma_{d}}$ plane for two sets of ($g_{d},\penalty 10000\ \sin\theta_{L}$).
3 Collider Analysis
In this section, we demonstrate the effect of the portal vector-like lepton on the Higgs production in association with a vector boson ($V=Z,\gamma_{d}$) at future muon collider. We first estimate the effect of pVLL on the production rate of $\mu^{+}\mu^{-}\to hV$ and compare this w.r.t. the SM prediction of $\mu^{+}\mu^{-}\to hZ$. The Feynman diagrams for $\mu^{+}\mu^{-}\to hV$ process are depicted in Fig. 2.
Figure 2: Feynman diagrams for associated production of Higgs and a vector boson ($V=Z,\gamma_{d}$) at muon collider ($\mu^{+}\mu^{-}\to hV$).
In order to generate the events and estimate the cross section we first implement the Lagrangian of Eq. (1) in Feynrules 2.0 [78]. The Universal Feynrules Output [79] of Feynrules is then interfaced with event generator MadGraph5_aMC@NLO (version $3.5.5$) [80] to estimate the leading order (LO) cross section in the context of muon collider.
We have chosen six benchmark points with parameter values consistent with the constraints mentioned in section 2.1, as shown in Table 2.
Benchmark Points $m_{{\mu_{{}_{p}}}}$ [TeV] $\sin{\theta_{L}}$ $g_{d}$ $\sin{\theta_{s}}$ $m_{{\gamma_{{}_{d}}}}$ [GeV]
BP1 11 $4.0\times 10^{-5}$ $0.45$ $0.05$ $90$
BP2 22 $2.0\times 10^{-5}$ $0.36$ $0.05$ $90$
BP3 33 $1.7\times 10^{-6}$ $0.50$ $0.05$ $90$
BP4 11 $4.0\times 10^{-5}$ $0.36$ $0.05$ $50$
BP5 22 $2.0\times 10^{-5}$ $0.30$ $0.05$ $50$
BP6 33 $1.7\times 10^{-6}$ $0.45$ $0.05$ $50$
Table 2: Representative benchmark points
3.1 $hZ$ vs $h\gamma_{d}$ at muon collider
Here we present a comparison between $\mu^{+}\mu^{-}\to h\gamma_{d}$ and $\mu^{+}\mu^{-}\to hZ$ production rates in the context of the model described in section 2 w.r.t. the SM $hZ$ production rate at future muon collider. In Fig. 3, we depict various ratios of cross sections defined in Eq. (23) for three different benchmark points.
$R_{1}$ $=$ $\frac{\sigma_{{}_{\rm BSM}}(\mu^{+}\mu^{-}\to hZ)}{\sigma_{{}_{\rm SM}}(\mu^{+}\mu^{-}\to hZ)}$ (23a)
$R_{2}$ $=$ $\frac{\sigma_{{}_{\rm unpol}}(\mu^{+}\mu^{-}\to h\gamma_{d})}{\sigma_{{}_{\rm SM}}(\mu^{+}\mu^{-}\to hZ)}$ (23b)
$R_{3}$ $=$ $\frac{\sigma^{\rm{s\mbox{-}ch}}_{{}_{\rm unpol}}(\mu^{+}\mu^{-}\to h\gamma_{d})}{\sigma_{{}_{\rm SM}}(\mu^{+}\mu^{-}\to hZ)}$ (23c)
$R_{4}$ $=$ $\frac{\sigma^{\rm{t\mbox{-}ch}}_{{}_{\rm unpol}}(\mu^{+}\mu^{-}\to h\gamma_{d})}{\sigma_{{}_{\rm SM}}(\mu^{+}\mu^{-}\to hZ)}$ (23d)
$R_{5}$ $=$ $\frac{\sigma^{\rm{t\mbox{-}ch}}_{{}_{\rm pol}}(\mu^{+}_{L}\mu^{-}_{R}\to h\gamma_{d})}{\sigma_{{}_{\rm SM}}(\mu^{+}\mu^{-}\to hZ)}$ (23e)
where,
$\sigma_{{}_{\rm SM}}(\mu^{+}\mu^{-}\to hZ)$ $=$ { $\begin{cases}\penalty 10000\ 1.36\penalty 10000\ {\rm fb},&{\rm for}\penalty 10000\ \sqrt{s}=3\penalty 10000\ {\rm TeV}\\ \penalty 10000\ 0.12\penalty 10000\ {\rm fb},&{\rm for}\penalty 10000\ \sqrt{s}=10\penalty 10000\ {\rm TeV}\end{cases}$ (24)
Figure 3: Ratios of cross sections ( $R_{i}$’s defined in Eq. (23) ) for different benchmark points at (a) $\sqrt{s}=3$ TeV and (b) $10$ TeV.
One can see that while the cross section for $\mu^{+}\mu^{-}\to hZ$ production in this scenario remains the same as that in the SM, a higher Higgs production rate is predicted in the $\mu^{+}\mu^{-}\to h\gamma_{d}$ channel. This is true even for a $\sin\theta_{L}$ as small [^4] Small $\sin{\theta_{L}}$ ($\sim 10^{-4}$ or less) is favoured by perturbative unitarity requirements for heavy muon mass beyond TeV scale as $\sim 10^{-5}$ with reasonable choices of other relevant parameters (see Table 2 for specification of benchmark points). The underlying reasons behind this is that $\mu_{R}^{\prime}$ and $\mu_{p_{{}_{R}}}^{\prime}$ share identical $SU(2)_{L}\times U(1)_{Y}$ charges. Hence, after transforming them into mass eigen-basis using Eq. (10) no off-diagonal $\mu^{-}_{R}-\mu^{-}_{p_{{}_{R}}}-Z$ interaction arises. However, the absence of $U(1)_{D}$ charge for the muon induces an off-diagonal $\mu_{R}^{-}-\mu^{-}_{p_{{}_{R}}}-\gamma_{d}$ interaction in the mass basis. It turns out that the corresponding interaction vertex $\mu^{+}_{L}-\mu_{p_{{}_{R}}}^{-}-Z$ is not present in this model while the $\mu^{+}_{L}-\mu_{p_{{}_{R}}}^{-}-\gamma_{d}$ (or $\mu^{-}_{R}-\mu_{p_{{}_{L}}}^{+}-\gamma_{d}$ ) vertex factor is proportional to $\sim g_{d}\sin\theta_{R}\cos\theta_{R}$ ( See Eq. (A.2) ). The mixing angle ($\sin\theta_{R}$) between the right chiral muon and heavy muon is given by
$\sin\theta_{R}=\dfrac{\sin\theta_{L}}{\sqrt{\dfrac{m_{\mu}^{2}}{m_{\mu_{p}}^{2}}\cos^{2}\theta_{L}+\sin^{2}\theta_{L}}}$ (25)
which can reach $\mathcal{O}(0.1)$ value for $m_{\mu_{{}_{p}}}$ in the TeV range even for a tiny $\sin\theta_{L}$ ($\sim 10^{-5}$). This is also independently validated by the result that the average unpolarized $t/u$-channel cross section for $\mu^{+}\mu^{-}\to h\gamma_{d}$ is dominated by the $\mu^{+}_{L}\mu^{-}_{R}\to h\gamma_{d}$ contribution and the contribution to the total unpolarized cross section predominantly comes from $t/u$-channel process compared to pure s-channel contribution. The later is suppressed due to the fact that $\mu^{+}_{L}-\mu^{-}_{R}-\gamma_{d}$ vertex is $\sim g_{d}\sin^{2}\theta_{R}$.
In Fig. 4 and Fig. 5, we plot $R_{2}$ which represents the ratio of average unpolarized cross section for $\mu^{+}\mu^{-}\to h\gamma_{d}$ to that of SM $\mu^{+}\mu^{-}\to hZ$ ( defined in Eq. (23) ) in the $m_{\gamma_{d}}$$m_{\mu_{p}}$ plane at $\sqrt{s}=3$ TeV and $\sqrt{s}=10$ TeV, respectively, to illustrate the effect of these two parameters on $R_{2}$ and hence the $h\gamma_{d}$ cross section for two different $g_{d}$ and $\sin\theta_{L}$ and fixed $\sin\theta_{s}$. The perturbative unitarity limit obtained using Eq. (15) is also displayed on the same plot. Since the requirement of perturbative unitarity disfavors low values of $v_{d}$, one cannot achieve an arbitrary high values of $R_{2}$ by going to lower values of dark photon mass or higher values of $g_{d}$ and a trade off is maintained in the choice of the two independent parameters $m_{\gamma_{d}}$ and $g_{d}$}.
Figure 4: Variation of ratio of cross sections ($R_{2}$) for the process $\mu^{+}\mu^{-}\to h\gamma_{d}$ with respect to the SM $hZ$ production rate in $m_{\gamma_{d}}$$m_{\mu_{{}_{p}}}$ plane assuming $\sqrt{s}=3$ TeV for two different set of ($g_{d},\penalty 10000\ \sin\theta_{L}$) and a fixed value of $\sin\theta_{s}$. (a) The left plot corresponds to $g_{d}=0.05,\penalty 10000\ \sin\theta_{s}=0.05,\penalty 10000\ \text{and}\penalty 10000\ \sin\theta_{L}=4\times 10^{-5}$ and (b) the right plot corresponds to $g_{d}=0.5,\penalty 10000\ \sin\theta_{s}=0.05,\penalty 10000\ \text{and}\penalty 10000\ \sin\theta_{L}=1.66\times 10^{-6}$. The region above the dashed (black) line is allowed by the constraint on $\omega_{f}$ described in Eq. (15).
Figure 5: Variation of ratio of cross sections ($R_{2}$) for the process $\mu^{+}\mu^{-}\to h\gamma_{d}$ with respect to the SM $hZ$ production rate in $m_{\gamma_{d}}$$m_{\mu_{{}_{p}}}$ plane assuming $\sqrt{s}=10$ TeV for two different set of ($g_{d},\penalty 10000\ \sin\theta_{L}$) and a fixed value of $\sin\theta_{s}$. (a) The left plot corresponds to $g_{d}=0.05,\penalty 10000\ \sin\theta_{s}=0.05,\penalty 10000\ \text{and}\penalty 10000\ \sin\theta_{L}=4\times 10^{-5}$ and (b) the right plot corresponds to $g_{d}=0.5,\penalty 10000\ \sin\theta_{s}=0.05,\penalty 10000\ \text{and}\penalty 10000\ \sin\theta_{L}=1.66\times 10^{-6}$. The region above the dashed (black) line is allowed by the constraint on $\omega_{f}$ described in Eq. (15).
Therefore, depending on the parameter space, Higgs production in association with a dark photon at muon collider can provide us a Higgs production rate higher than $hZ$ production by a factor ranging from $1$ to $100$.
3.2 Analysis of $h\gamma_{d}$ process
In the remaining part of this article we focus on the $h\gamma_{d}$ production at the future muon collider. As discussed above, the rate of this process can be significantly large compared to the $hZ$ production in a wide range of dark photon mass. Hence, it can contribute as an additional channel to the Higgs production at muon collider.
In Fig. 6 we present the LO cross sections for $\mu^{+}\mu^{-}\to h\gamma_{d}$ as a function of the center-of-mass energy.
Figure 6: LO cross section for $\mu^{+}\mu^{-}\to h\gamma_{d}$ as a function of the muon collider center-of-mass energy for various choices of the $m_{\mu_{p}}$ and $\sin\theta_{L}$ assuming (a) $\sin\theta_{s}=0.05,\penalty 10000\ g_{d}=0.36,\penalty 10000\ m_{\gamma_{d}}=50\penalty 10000\ \text{GeV}$ and (b) $\sin\theta_{s}=0.05,\penalty 10000\ g_{d}=0.50,\penalty 10000\ m_{\gamma_{d}}=90\penalty 10000\ \text{GeV}$.
The LO cross sections for the process $\mu^{+}\mu^{-}\to h\gamma_{d}$ for various benchmark points mentioned in Table 2 at the muon collider center-of-mass energies $\sqrt{s}=3$ TeV and $10$ TeV are presented in Table 3.
$\sqrt{s}$ $\sigma(\mu^{+}\mu^{-}\to h\gamma_{d})$ [fb]
[TeV] BP1 BP2 BP3 BP4 BP5 BP6
3 $11.1$ $22.8$ $0.064$ $47.2$ $114$ $0.436$
10 $2.66$ $10.9$ $0.056$ $11.4$ $55.0$ $0.388$
Table 3: LO cross sections for $\mu^{+}\mu^{-}\to h\gamma_{d}$ at $\sqrt{s}=3$ TeV and $10$ TeV for various choices of our benchmark points
The cross section for $h\gamma_{d}$ production strongly depends on dark photon mass and dark gauge coupling as well as on other model parameters (see Fig. 4, Fig. 5, as well as the corresponding vertex factors in Appendix A and the squared matrix elements in Appendix B). The cross section for the $\mu^{+}\mu^{-}\to h\gamma_{d}$ process as a function of the dark gauge coupling ($g_{d}$) can be parametrized as
$\sigma_{{}_{\rm tot}}(\mu^{+}\mu^{-}\to h\gamma_{d})$ $\approx$ $g_{d}^{2}(c_{1}\frac{g_{d}}{m_{\gamma_{d}}}+c_{2})^{2}\mathcal{A}_{1}\penalty 10000\ +\penalty 10000\ m_{\gamma_{d}}g_{d}^{3}(c_{1}\frac{g_{d}}{m_{\gamma_{d}}}+c_{2})\mathcal{A}_{2}$ (26)
where $c_{1},\penalty 10000\ c_{2}$ and $\mathcal{A}_{1},\penalty 10000\ \mathcal{A}_{2}$ are in general functions of ($m_{\mu_{p}},\penalty 10000\ \sin\theta_{L},\penalty 10000\ \sin\theta_{s}$) and ($\sqrt{s}$, $m_{\gamma_{d}}$, $m_{\mu_{p}}$, $\sin\theta_{L}$, $\sin\theta_{s}$), respectively. For example, assuming $m_{\gamma_{d}}=90\penalty 10000\ \text{GeV},\penalty 10000\ m_{\mu_{p}}=1\penalty 10000\ \text{TeV},\penalty 10000\ \sin\theta_{L}=4\times 10^{-5}\penalty 10000\ \text{and}\penalty 10000\ \sin\theta_{s}=0.05$ one has, at $g_{d}=0.45$ and $\sqrt{s}=3$ TeV
$c_{1}=16.56\penalty 10000\ \text{GeV},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ c_{2}=1.72\times 10^{-8}$
$\mathcal{A}_{1}=7.96\times 10^{3}\penalty 10000\ \text{fb},\penalty 10000\ \penalty 10000\ \mathcal{A}_{2}=0.145\penalty 10000\ \text{fb\penalty 10000\ GeV}^{-1}$
While writing down the above parametric equation a particular helicity configuration of the initial state muons have been considered that gives dominant contribution to the total unpolarized cross section as explained earlier. Also we have ignored the pure $s$-channel and any $t/u$-channel muon mediated contributions. In Eq. (26), the first term corresponds to pure t/u-channel heavy muon mediated contribution while the second term represents its interference with the $s$-channel diagram.
The dark photon in this scenario is considered to be invisible and hence gives rise to missing enrgy signature at the muon collider. The Higgs on the other hand predominantly decays to a pair of bottom quarks. Therefore, the final state consists of $b\bar{b}$ along with missing energy at muon collider (see Fig. 7).
Figure 7: Feynman diagrams for the subprocess, $\mu^{+}\mu^{-}\to h(b\bar{b})+\gamma_{d}$ leading to $b\bar{b}$ plus missing energy final state.
The dominant SM backgrounds that contribute to the $b\bar{b}$ plus missing energy final state, namely, $\mu^{+}\mu^{-}\to b\bar{b}\nu\bar{\nu}$ process comes from $hZ,\penalty 10000\ ZZ,\penalty 10000\ h\nu\bar{\nu},\penalty 10000\ Z\nu\bar{\nu}$ and non-resonant $b\bar{b}\nu\bar{\nu}$ productions at muon collider. They can be sub-divided into $2\to 2$, $2\to 3$ and $2\to 4$ categories as follows
  • 0 $\mu^{+}\mu^{-}\to hZ\to(b\bar{b})(\nu\bar{\nu})$, \__color_backend_reset:0 () \__color_backend_reset:    0 $\mu^{+}\mu^{-}\to ZZ\to(b\bar{b})(\nu\bar{\nu})$, \__color_backend_reset:0 () \__color_backend_reset:    0 $\mu^{+}\mu^{-}\to h\nu\bar{\nu}\to(b\bar{b})\penalty 10000\ \nu\bar{\nu}$, \__color_backend_reset:0 () \__color_backend_reset:
    0 $\mu^{+}\mu^{-}\to Z\nu\bar{\nu}\to(b\bar{b})\penalty 10000\ \nu\bar{\nu}$, \__color_backend_reset:0 () \__color_backend_reset:    0 $\mu^{+}\mu^{-}\to b\bar{b}\penalty 10000\ \nu\bar{\nu}$ (non-resonant). \__color_backend_reset:0 () \__color_backend_reset:
Some representative Feynman diagrams corresponding to these SM background subprocesses are shown in Fig. 8.
Figure 8: Feynman diagrams for various SM background subprocesses leading to $b\bar{b}\nu\bar{\nu}$ final state in $\mu^{+}\mu^{-}$ collisions.
In Table 4, we tabulate the cross sections for the aforementioned SM background subprocesses at $3$ and $10$ TeV $\mu^{+}\mu^{-}$ collision energies. It is obvious that the dominant contribution to the total SM background comes from the resonant $b\bar{b}$ production.
Background $\sigma$ [fb] $\sigma$ [fb]
processes at $\sqrt{s}=3$ TeV at $\sqrt{s}=10$ TeV
$\mu^{+}\mu^{-}\to hZ\to(b\bar{b})(\nu\bar{\nu})$ $0.18$ $0.02$
$\mu^{+}\mu^{-}\to ZZ\to(b\bar{b})(\nu\bar{\nu})$ $1.45$ $0.18$
$\mu^{+}\mu^{-}\to h\nu\bar{\nu}\to(b\bar{b})\nu\bar{\nu}$ $318$ $540$
$\mu^{+}\mu^{-}\to Z\nu\bar{\nu}\to(b\bar{b})\nu\bar{\nu}$ [^5] The cross section quoted in the table for this process includes resonant $\nu\bar{\nu}$ contribution from $\mu^{+}\mu^{-}\to ZZ$ as well $305$ $526$
$\bm{\mu^{+}\mu^{-}\to b\bar{b}\nu\bar{\nu}}\penalty 10000\ ({\rm{\bf total}})$ $\bm{641}$ $\bm{1101}$
Table 4: LO cross sections (in fb) for various SM background subprocesses contributing to the $b\bar{b}\nu\bar{\nu}$ final state in $\mu^{+}\mu^{-}$ collision at $\sqrt{s}=3$ TeV and $10$ TeV. The values in the last row (boldface) take into account contributions from all possible subprocesses leading to $b\bar{b}\nu\bar{\nu}$ final state including those mentioned in the table.
Events generated using the event generator MadGraph5_aMC@NLO for both the signal and the SM background processes are then interfaced with Pythia8 (version 8.3) [81] for showering, hadronization, decay and further collider analysis. The detector response in the context of future muon collider is parametrized using a Gaussian smearing function as given in Eq. (27) following the Ref. [82, 83].
$\frac{\sigma_{{}_{E}}}{E}$ $=$ $\sqrt{\left(\frac{N}{E}\right)^{2}+\left(\frac{S}{\sqrt{E}}\right)^{2}+C^{2}}$ (27)
where $E$ is the energy associated with the final state object and $\sigma_{{}_{E}}$ is the standard deviation in the measurement of $E$. The parameters $N,\penalty 10000\ S$ and $C$ represents respectively the noise, stochastic and the constant terms. Depending on the detector these parameters have different values for specific final state objects. In the context of future muon collider for jets in the central region ($|\eta|<2.5$), we have used the following set of values for these parameters [84]
$N=0,\penalty 10000\ \penalty 10000\ S=0.3,\penalty 10000\ \penalty 10000\ C=0.05$
In Fig. 9, we plot the central-jet [^6] Jets with transverse momentum $p_{{}_{T}}>20$ GeV and pseudo-rapidity $|\eta|<2.5$ number distributions for various signal benchmark points. We have used anti-kt jet clustering algorithm [85] with clustering radius R=0.4 using FastJet (version 3.5.0) [86]. This has been done keeping in mind the fact that at the muon collider center-of-mass energies ($\sqrt{s}=3\penalty 10000\ \text{TeV}\penalty 10000\ \text{and}\penalty 10000\ 10\penalty 10000\ \text{TeV})$, both the Higgs and the dark photon are always highly boosted, each sharing almost half of the available center-of-mass energies. Therefore, most of the signal events are expected to have one collimated $b\bar{b}$ pair originating from the Higgs decays resulting into a single narrow jet. Only a fraction of signal events due to final state radiation, showering and hadronization effect result into events with more than one central-jet.
It turns out at 3 TeV center-of-mass energy a large fraction ($\sim 87-90$%) of the signal events consists of exactly one central-jet and remaining 101310-13% corresponds to exactly two central-jet events. This pattern changes at 10 TeV center-of-mass energy. In this case, almost all the signal events ($\gtrsim 96$%) correspond to one central-jet, as expected. The central-jet number distribution corresponding to the total SM background in the $b\bar{b}\nu\bar{\nu}$ final state is also displayed on the same figure. One can see that the SM background events are distributed among $0,1,2\penalty 10000\ \text{or}\penalty 10000\ 3$ central-jet events categories. At $\sqrt{s}=3$ TeV the distribution has a peak around Nc-jet = 2 containing almost 5454% of the total background events. At 10 TeV center-of-mass energy, there is a significant decrease (increase) in number of events in exactly two (zero) central-jet event categories.
Figure 9: Central-jet (Nc-jet) multiplicity distribution at $\sqrt{s}=3$ TeV and $10$ TeV for three different signal benchmark points and the total SM background.
Figure 10: Central-jet (Nc-jet) multiplicity distribution at $\sqrt{s}=3$ TeV and $10$ TeV for various SM background subprocesses and the total background.
This behavior arises because a significant fraction of the background originates from $2\to 3$ processes such as $h\nu\bar{\nu}$ and $Z\nu\bar{\nu}$ [^7] We have not explicitly shown corresponding distributions for $Z\nu\bar{\nu}$ and also in the context of distributions of various kinematical variables later on as it will also contain the contributions coming from $\mu^{+}\mu^{-}\to ZZ$ process.. In these cases, the Higgs or $Z$ boson produced in association with a neutrino pair is generally not boosted, often resulting in multi-jet final states. Similar arguments apply to the continuum $b\bar{b}\nu\bar{\nu}$ background. In contrast, the SM $hZ$ and $ZZ$ backgrounds, although both are $2\to 2$ processes, display different central-jet number distributions. This distinction arises because the $hZ$ is predominantly an $s$-channel process, whereas the $ZZ$ process proceeds via $t/u$-channel muon mediation, leading to different angular distributions of the final-state particles. Consequently, the $ZZ$ process is expected to give more zero central-jet events than $hZ$. This feature is also reflected in Fig. 10.
This motivates us to focus on two different final states, namely, leptonically quiet exactly one central-jet and exactly two central-jets, both in conjunction with missing energy at two different $\mu^{+}\mu^{-}$ collision energies ($\sqrt{s}=3$ TeV and $10$ TeV).
Figure 11: Transverse momentum ($p_{{}_{T}}$) distributions of the (a) jet and (b) the di-jet in the Nc-jet = 1 and Nc-jet = 2 final states, respectively, at $3$ TeV muon collider center-of-mass energy for both signal and the SM background events.
Figure 12: Transverse momentum ($p_{{}_{T}}$) distributions of the (a) jet and (b) the di-jet in the Nc-jet = 1 and Nc-jet = 2 final states, respectively, at $10$ TeV muon collider center-of-mass energy for both signal and the SM background events.
As already mentioned, Higgs boson produced at the future muon collider center-of-mass energies in a $2\to 2$ process are typically highly boosted. The Higgs subsequently decays into a pair of bottom quarks which due to their small angular separation, manifest as a single collimated jet in the detector. In Fig. 11 and 12, we plot the transverse momentum distributions of a jet or di-jet system, as applicable, for two different muon collider center-of-mass energies. The purpose of these distributions are two-fold. They provide an explanation of Fig. 9 and 10 and also give a motivation for the following discussion.
To identify and analyze the boosted Higgs jets, we employ jet substructure analysis and dedicated Higgs tagging techniques. For this purpose we have used MassDropTagger [87] to identify a boosted Higgs structure inside jets. These jets are obtained after re-clustering all the hadrons in a particular final state with Cambridge-Aachen (CA) jet clustering algorithm [88] and energy-scheme recombination assuming jet-clustering radius [^8] We call these jets as CA1.2 jets. Also the above analysis is not too sensitive to the choice of the jet clustering radius values. The algorithm works fine in the range $0.8\leq R\leq 1.4$. $R=1.2$. The CA algorithm is chosen because it clusters particles based solely on angular distance, making it well-suited for subsequent substructure analyses.
Below we briefly summarize the algorithm behind the jet substructure analysis implemented in MassDropTagger:
  1. 1.
    The leading-$p_{{}_{T}}$ jet (with transverse momentum $p_{{}_{T}}>200\penalty 10000\ \mathrm{GeV}$) is first passed through the MassDropTagger procedure. The algorithm iteratively de-clusters the jet to find an imprint of a heavy-particle by applying the following two conditions characterized by the dimensionless parameters $\mu$ and $y_{\text{cut}}$:
    • i)
      Mass-drop condition : $\frac{m_{j_{1}}}{m_{j}}<\mu$
    • ii)
      $\text{Symmetry condition:}\penalty 10000\ y=\frac{\min\!\left(p_{Tj_{1}}^{2},\,p_{Tj_{2}}^{2}\right)}{m_{j}^{2}}\,\Delta R_{j_{1}j_{2}}^{2}\;>\;y_{\text{cut}}\,.$
    where $j_{1}$ and $j_{2}$ (with $m_{j_{1}}>m_{j_{2}}$) denote the two sub-jets obtained by undoing the final stage of clustering of a parent jet $j$. Here, $\Delta R_{j_{1}j_{2}}$ (with $\Delta R=\sqrt{\Delta\eta^{2}+\Delta\phi^{2}}$) is the angular separation between the two sub-jets in the $\eta-\phi$ plane. The values of the dimensionless parameters $\mu$ and the $y_{\rm cut}$ are $0.67$ and $0.09$, respectively [87].
  2. 2.
    If both the above criteria are satisfied, the jet is tagged as a boosted heavy-particle candidate consisting of two sub-jets $j_{1}$ and $j_{2}$. Otherwise, the lighter sub-jet is discarded, and the procedure is repeated by treating $j_{1}$ as the parent jet until a valid substructure is identified. We select an event for further analysis if it contains at least one ${\rm CA}_{1.2}$ jet with a substructure.
As discussed earlier in this section, the dominant contribution to the total SM background cross section (approximately 0.641 pb at $\sqrt{s}=3$ TeV) arises from $2\to 3$ processes such as $\mu^{+}\mu^{-}\to h\nu\bar{\nu}$ and $\mu^{+}\mu^{-}\to Z\nu\bar{\nu}$. The application of jet-substructure techniques or the use of the transverse momentum of the $b\bar{b}$ system, in combination with missing energy, proves highly effective in suppressing these dominant SM background components. Consequently, only the $2\to 2$ processes like $\mu^{+}\mu^{-}\to ZZ$ and $\mu^{+}\mu^{-}\to hZ$ surpass the substructure requirement. However, the cross sections for these later processes are significantly smaller compared to the total SM background cross section (see Table 4).
3.3 Kinematic variables
In this subsection we discuss few useful kinematic variables and choice of selection cuts. The next stage of our cut based analysis utilizes both the missing energy and invariant mass of a jet (or di-jet system) to reduce/eliminate the SM backgrounds.
  • Missing energy ($\not${E}): In collider experiments, missing energy corresponds to the total energy carried by the invisible particles which remain unobserved at the detector level. In a particular final state consisting of both visible and invisible particles, using the energy conservation we can find the missing energy of the system as
[…]
    where, $\sqrt{s}$ is center-of-mass energy of the collisions and $E_{\text{vis}}$ is the total energy of all the visible final state particles.
    In collider investigations, the visible particles in general consist of the hadrons clustered as jets, isolated photons and leptons, and soft unclustered components. Thus, the expression of missing energy at the detector level can be written as
[…]
    In Fig. 13, we plot the missing energy distributions for the signal events for various benchmark scenarios corresponding to two different final states at center-of-mass energy $\sqrt{s}=3\penalty 10000\ \text{TeV}$. The corresponding distributions at 10 TeV center-of-mass energy are shown in Fig. 14. The contribution to the missing energy for the signal events is due to the invisible dark photon and the distribution is expected to peak $\sim{\sqrt{\rm s}/2}$ (neglecting the finite mass effect) in a $2\to 2$ process. However, due to the presence of additional neutrinos coming from $B$ meson decays the missing energy distribution extends beyond $\sqrt{\rm s}/2$}. The SM background events in the $b\bar{b}\nu\bar{\nu}$ final state has dominant contribution from the $2\to 3$ processes such as $h\nu\bar{\nu}$ and $Z\nu\bar{\nu}$. Together they form the 99 % of the total cross section and only a very small fraction of it comes from $2\to 2$ processes like $ZZ$ and $hZ$. The pair of neutrinos in the $h\nu\bar{\nu}$ or $Z\nu\bar{\nu}$ processes carries maximum share of the available center-of-mass energy. Hence, the missing energy distribution for the SM background events has a peak mostly around $\sqrt{s}$. This is reflected in the missing energy distributions for individual SM background components and the total SM background as depicted in the same figures. Therefore, missing energy variable is an extremely useful discriminator between the signal and the background.
    Figure 13: Missing energy ($\not${E}) distributions at $\sqrt{s}=3$ TeV in the (a) Nc-jet = 1 and (b) Nc-jet = 2 final states for both signal and SM background events.
    Figure 14: Missing energy ($\not${E}) distributions at $\sqrt{s}=10$ TeV in the (a) Nc-jet = 1 and (b) Nc-jet = 2 final states for both signal and SM background events.
  • $\bm{b\bar{b}}$ invariant mass ($M_{b\bar{b}}$) : In our analysis the invariant mass associated with the $b\bar{b}$ system is extracted in two different ways using the following formula
    $M_{12}^{2}=(p_{1}+p_{2})^{2}=(E_{1}+E_{2})^{2}-|\vec{p_{1}}+\vec{p_{2}}|^{2}$ (30)
    where $p_{1}=(E_{1},|\vec{p_{1}}|)$ and $p_{2}=(E_{2},|\vec{p_{2}}|)$ are the four momentum associated with the two sub-jets (jets) in the final state comprising one (two) central-jet. In events with exactly one central-jet we combine the four momentum of the two sub-jets inside a boosted $\rm{CA}_{1.2}$ jet to get the invariant mass distribution. Whereas in the Nc-jet = 2, case we combine the four momentum of two AK4 jets in order to get the invariant mass associated with the di-jet system. The corresponding distributions in these two final states are plotted in Fig. 15 and 16. One can see that for signal events this distribution has a peak near the Higgs mass. Whereas the same for the background components peak at varying locations, namely, at the mass of the Higgs and $Z$ boson as expected. One also sees a peak at a very low invariant mass which has origin in the final state radiation and hadronization effect from one of the bottom quarks only, and may give rise to jet substructure. In events with two central-jets in final state one has two clear peaks at the values of $Z$ and Higgs mass in the invariant mass distribution corresponding to the background subprocesses. This is due to the presence of $ZZ,\penalty 10000\ Z\nu\bar{\nu}$ component as well as $hZ$ and $h\nu\bar{\nu}$ components of the total SM background. The tail in all these invariant mass distributions at values lower than the peak values is due to the missing momentum component associated with the invisible neutrinos produced in $B$ decays.
Figure 15: Invariant mass ($M_{b\bar{b}}$) distribution at $\sqrt{s}=3$ TeV in the (a) Nc-jet = 1 and (b) Nc-jet = 2 final state for both signal and background events.
Figure 16: Invariant mass ($M_{b\bar{b}}$) distribution at $\sqrt{s}=10$ TeV in the (a) Nc-jet = 1 and (b) Nc-jet = 2 final state for both signal and background events.
One can see that the normalized signal distributions are mostly insensitive to the intermediate heavy muon mass. While presenting the kinematic distributions for the signal benchmark points (BPs) we have chosen (BP1, BP2 and BP3) as our representative BPs corresponding to three different heavy muon masses ($1,\penalty 10000\ 2\penalty 10000\ \text{and}\penalty 10000\ 3$ TeV). Similar distributions exist for other signal benchmark points {BP4–BP6} corresponding to $m_{\gamma_{d}}=50\penalty 10000\ \text{GeV}$ as well. However, the effect of the dark photon mass on the normalized distributions are not significant enough. Hence, we choose not to display them here as they closely resemble the distributions already presented.
3.4 Cuts
To separate the signal from the corresponding SM background, we have used the following selection cuts on the observables associated with the individual final states.
[…]
  • Jet substructure requirement
  • $M_{b\bar{b}}>M_{0}$, where $M_{0}=95$ GeV.
4 Results
[…]
BPs Production Cross sections (fb): Nc-jet = 1 final state Cross sections (fb): Nc-jet = 2 final state
cross-section Basic $\not${E} cut Jet substructure $M_{b\bar{b}}$ cut Basic $\not${E} cut Jet substructure $M_{b\bar{b}}$ cut
[…]
BP1 $11.1$ $9.70$ $9.51$ $9.39$ $7.36$ $1.42$ $1.38$ $1.30$ $1.24$
BP2 $22.8$ $20.7$ $20.3$ $20.1$ $15.7$ $2.03$ $1.97$ $1.84$ $1.76$
BP3 $0.064$ $0.059$ $0.058$ $0.057$ $0.044$ $0.005$ $0.005$ $0.005$ $0.005$
BP4 $47.2$ $41.1$ $40.3$ $39.8$ $31.3$ $5.97$ $5.79$ $5.49$ $5.24$
BP5 $114$ $104$ $102$ $101$ $78.8$ $9.95$ $9.63$ $8.97$ $8.65$
BP6 $0.436$ $0.401$ $0.393$ $0.388$ $0.303$ $0.035$ $0.034$ $0.031$ $0.030$
Table 5: The effect of cut-flow on the signal cross section for various representative benchmark points in two different final states ( Nc-jet = 1 and Nc-jet = 2) at $\sqrt{s}=3$ TeV assuming BR($h\to b\bar{b}$) =1=1.
Background Production Cross sections (fb): Nc-jet = 1 final state Cross sections (fb): Nc-jet = 2 final state
processes cross-section Basic $\not${E} cut Jet substructure $M_{b\bar{b}}$ cut Basic $\not${E} cut Jet substructure $M_{b\bar{b}}$ cut
[…]
$b\bar{b}\nu\bar{\nu}\penalty 10000\ (hZ)$ $0.178$ $0.164$ $0.161$ $0.160$ $0.124$ $0.014$ $0.013$ $0.012$ $0.012$
$b\bar{b}\nu\bar{\nu}\penalty 10000\ (ZZ)$ $1.45$ $0.495$ $0.473$ $0.368$ $0.014$ $0.127$ $0.124$ $0.120$ $0.007$
$b\bar{b}\nu\bar{\nu}\penalty 10000\ (h\nu\bar{\nu})$ $318$ $95.5$ $4.50$ $0.614$ $0.347$ $187$ $1.42$ $1.03$ $0.897$
$\bm{b\bar{b}\nu\bar{\nu}}$ (tot) $\bm{641}$ $\bm{222}$ $\bm{10.9}$ $\bm{4.00}$ $\bm{0.538}$ $\bm{351}$ $\bm{4.75}$ $\bm{3.97}$ $\bm{1.44}$
Table 6: The effect of cut-flow on the cross-sections for various SM background subprocesses and the total background in two different final states ( Nc-jet = 1 and Nc-jet = 2) at $\sqrt{s}=3$ TeV.
BPs Production Cross sections (fb): Nc-jet = 1 final state Cross sections (fb): Nc-jet = 2 final state
cross-section Basic $\not${E} cut Jet substructure $M_{b\bar{b}}$ cut Basic $\not${E} cut Jet substructure $M_{b\bar{b}}$ cut
[…]
BP1 $2.66$ $2.55$ $2.53$ $2.49$ $1.84$ $0.027$ $0.026$ $0.024$ $0.023$
BP2 $10.9$ $10.8$ $10.7$ $10.5$ $7.72$ $0.040$ $0.039$ $0.036$ $0.035$
BP3 $0.056$ $0.056$ $0.056$ $0.055$ $0.040$ $\sim 10^{-4}$ $\sim 10^{-4}$ $\sim 10^{-4}$ $\sim 10^{-4}$
BP4 $11.4$ $10.9$ $10.8$ $10.6$ $7.85$ $0.114$ $0.112$ $0.102$ $0.099$
BP5 $55.0$ $54.5$ $54.0$ $53.1$ $38.9$ $0.207$ $0.203$ $0.184$ $0.180$
BP6 $0.388$ $0.386$ $0.383$ $0.376$ $0.276$ $\sim 6\cdot 10^{-4}$ $\sim 6\cdot 10^{-4}$ $\sim 6\cdot 10^{-4}$ $\sim 6\cdot 10^{-4}$
Table 7: The effect of cut-flow on the signal cross section for various representative benchmark points in two different final states ( Nc-jet = 1 and Nc-jet = 2) at $\sqrt{s}=10$ TeV assuming BR($h\to b\bar{b}$) =1=1.
Background Production Cross sections (fb): Nc-jet = 1 final state Cross sections (fb): Nc-jet = 2 final state
processes cross-section Basic $\not${E} cut Jet substructure $M_{b\bar{b}}$ cut Basic $\not${E} cut Jet substructure $M_{b\bar{b}}$ cut
[…]
$b\bar{b}\nu\bar{\nu}\penalty 10000\ (hZ)$ $0.016$ $0.016$ $0.016$ $0.015$ $0.011$ $10^{-5} <$$10^{-5} <$$10^{-5} <$$10^{-5}$
$b\bar{b}\nu\bar{\nu}\penalty 10000\ (ZZ)$ $0.183$ $0.045$ $0.044$ $0.043$ $0.002$ $\sim 2\cdot 10^{-4}$ $\sim 2\cdot 10^{-4}$ $\sim 2\cdot 10^{-4}$ $10^{-4}$
$b\bar{b}\nu\bar{\nu}\penalty 10000\ (h\nu\bar{\nu})$ $540$ $167$ $0.717$ $0.588$ $0.451$ $246$ $0.079$ $0.077$ $0.073$
$\bm{b\bar{b}\nu\bar{\nu}}$ (tot) $\bm{1101}$ $\bm{373}$ $\bm{4.03}$ $\bm{3.53}$ $\bm{0.673}$ $\bm{466}$ $\bm{0.488}$ $\bm{0.427}$ $\bm{0.310}$
Table 8: The effect of cut-flow on the cross-sections for various SM background subprocesses and the total background in two different final states ( Nc-jet = 1 and Nc-jet = 2) at $\sqrt{s}=10$ TeV.
It is clear from the Tables 5 and 7 that at muon collider center-of-mass energies the signal events dominantly contribute in the exactly one central-jet final state than the other one because of the fact that it is a $2\to 2$ process. Moreover, the contribution of signal events in exactly one central-jet final state increases with increasing mass of the heavy particle (VLL) in the $t/u$ channel. This is evident if we compare the event rate after basic selection criteria for various benchmark points ({BP1-BP3} or {BP4-BP6}) which corresponds to pVLL mass in the range {1-3} TeV. This feature is also visible in Fig. 9.
For the background events coming from the $\mu^{+}\mu^{-}\to hZ$ subprocess more than 9292% of the events contribute to Nc-jet = 1 final state and the effects of cut-flow on these events are similar to that of the signal events. The contribution of $\mu^{+}\mu^{-}\to ZZ$ subprocess to the Nc-jet = 1 final state is 3434% (2525%) of its production cross section and a very negligible fraction contributes in the Nc-jet = 2 final state at $\sqrt{s}=3\penalty 10000\ (10)$ TeV (see Table 6 and 8). This is because a large fraction of events coming from this subprocess contributes in zero central-jet events, as can be seen in Fig. 10. On the other hand, though $h\nu\bar{\nu}$ (and also $Z\nu\bar{\nu}$) contributes dominantly to Nc-jet = 2 final state, however, its contribution after applying the selection criteria mentioned earlier can be reduced substantially.
The effect of our selection criteria on the signal events is less severe in both the final states as can be seen from the cut-flow tables. Whereas one has a factor of $\sim\frac{1}{400}$ and $\frac{1}{550}$ suppression in the SM background rate in the exactly one central-jet final state after applying the combination of cuts mentioned in section 3.4 at $\sqrt{s}=3$ TeV and $10$ TeV, respectively. This factor for the exactly two central-jet final state is $\sim\frac{1}{250}$ for $\sqrt{s}=3$ TeV and even larger at 10 TeV center-of-mass energy. The significant effect on the background comes from the missing energy cut as already explained in section 3.
The cut-flow tables corresponding to the signal events also suggest that the effect of the dark photon mass on the kinematics, in particular, on the cut efficiencies is very small as long as $m_{\gamma_{d}}\ll\sqrt{s}$. However, it plays a significant role in the estimation of the cross section as it depends on values the of dark photon mass as well as the dark gauge coupling.
In Table 9, we quote the signal significances for $3$ TeV and $10$ TeV muon collider center-of-mass energies for various choices of our benchmark points for two different final states assuming an integrated luminosity of $100$ fb-1. To estimate the signal significance ($\sigma$) and set a $2\sigma$ exclusion limit we have used the following formula [89, 90, 91]
$\sigma\;=\;\sqrt{\,2\left(S-B\ln\!\left(1+\frac{S}{B}\right)\right)}.$ (31)
where, $S$ is number of signal events and $B$ is that for the background at a given integrated luminosity. While estimating the significance we have also taken into account the appropriate branching ratio ($57\%$) for $h\to b\bar{b}$ decay to get the actual number of signal events.
Signal significance ($\sigma$)
BPs $\sqrt{s}=3$ TeV $\sqrt{s}=10$ TeV
Nc-jet = 1 Nc-jet = 2 Nc-jet = 1 Nc-jet = 2
BP1 $24.6$ $8.34$ $9.12$ $0.23$
BP2 $38.5$ $11.1$ $24.6$ $0.35$
BP3 $0.34$ $0.04$ $0.27$ -
BP4 $56.4$ $16.6$ $24.9$ $0.96$
BP5 $92.2$ $23.6$ $62.9$ $1.67$
BP6 $2.14$ $0.14$ $1.78$ -
Table 9: Estimated signal significance in the context of $\mu^{+}\mu^{-}\to h\gamma_{d}$ process in two different final states ( Nc-jet = 1 and Nc-jet = 2) under consideration at two different muon collider center-of-mass energies ($\sqrt{s}=3$ and $10$ TeV) assuming an integrated luminosity of 100 fb-1.
The increase in significance for BP2 compared to BP1 is due to higher signal cross section at this benchmark point even though the mass of the pVLL in the $t/u$-channel propagator is higher. Also one has a better estimate of the signal significance for the set of BPs corresponding to dark photon mass 50 GeV compared to that for dark photon mass 90 GeV as the cross section for $h\gamma_{d}$ production is expected to increase with decreasing dark photon mass.
In Fig. 17, we present the $2\sigma$ exclusion limit in the $m_{\gamma_{d}}-m_{\mu_{p}}$ plane at $\sqrt{s}=3$ TeV assuming an integrated luminosity of 1 ab-1 in two different final states for two sets of $g_{d}$ and $\sin\theta_{L}$ keeping $\sin\theta_{s}$ fixed at $0.05$. The same at $\sqrt{s}=10$ TeV and an integrated luminosity of $10$ ab-1 is presented in Fig. 18. The Fig. 19 represents the $2\sigma$ exclusion limit in the $g_{d}-m_{\gamma_{d}}$ plane corresponding to $\sqrt{s}=3$ TeV and an integrated luminosity of $1$ ab-1. The regions allowed by the constraint presented in Eq. (15) and the current muon g2g-2 measurements at $1\sigma$ level are represented by the one above the dashed (black) and the solid (black) lines, respectively, in all these figures.
Figure 17: $2\sigma$ exclusion limit in the $m_{\gamma_{d}}-m_{\mu_{p}}$ plane for two different final states at $3$ TeV muon collider center-of-mass energy for (a) $g_{d}=0.05,\penalty 10000\ \sin\theta_{L}=4\times 10^{-5},\penalty 10000\ \sin\theta_{s}=0.05$ and (b) $g_{d}=0.5,\penalty 10000\ \sin\theta_{L}=1.66\times 10^{-6},\penalty 10000\ \sin\theta_{s}=0.05$ assuming an integrated luminosity of 1 ab-1. The exclusion limits in Nc-jet = 1 (pink) and Nc-jet = 2 (blue) final states are represented by the region below the dash-dot lines. The region above the dashed (black) line is allowed by the constraint represented in Eq. (15) while the muon g2g-2 allowed region is represented by the one above solid (black) line.
Figure 18: $2\sigma$ exclusion limit in the $m_{\gamma_{d}}-m_{\mu_{p}}$ plane in Nc-jet = 1 final state at $10$ TeV muon collider center-of-mass energy for (a) $g_{d}=0.05,\penalty 10000\ \sin\theta_{L}=4\times 10^{-5},\penalty 10000\ \sin\theta_{s}=0.05$ and (b) $g_{d}=0.5,\penalty 10000\ \sin\theta_{L}=1.66\times 10^{-6},\penalty 10000\ \sin\theta_{s}=0.05$ assuming an integrated luminosity of 10 ab-1. The $2\sigma$ exclusion limit is represented by dash-dot line (pink). The region above the dashed (black) line is allowed by the constraint represented in Eq. (15) while the muon g2g-2 allowed region is represented by the one above solid (black) line.
Figure 19: $2\sigma$ exclusion limit (represented by dashed-dot line) in the $g_{d}-m_{\gamma_{d}}$ plane using the data corresponding to $\sqrt{s}=3$ TeV and an integrated luminosity of $1$ ab-1 in the Nc-jet = 1 (pink) and Nc-jet = 2 (blue) final states for two different sets of $\sin\theta_{L}$ and $m_{\mu_{p}}:$ (a) $\sin\theta_{L}=4\times 10^{-5}$, $m_{\mu_{p}}=1$ TeV and (b) $\sin\theta_{L}=1.66\times 10^{-6}$, $m_{\mu_{p}}=3$ TeV. The region above the dashed (black) line is allowed by the constraint represented in Eq. (15) while the muon g2g-2 allowed region is represented by the one above solid (black) line.
5 Conclusion
Motivated by the prospects of Higgs physics at proposed future lepton colliders and precise measurements of the Higgs properties, we have investigated the implications of a portal vector-like lepton on the Higgs production in association with a neutral gauge boson ($Z/\gamma_{d}$) at the proposed future muon collider experiment and its consequence on dark photon. In particular, we focus on the Higgs and dark photon associated production via $t$ and $u$ channel pVLL exchange in addition to the dark photon mediated $s$-channel contribution. To avoid constraints coming from the low energy flavor changing observables we consider pVLL mixing only with one generation of lepton family, in particular, the second generation charged lepton. We also present a detailed comparison of the $h\gamma_{d}$ production rate with that of the $hZ$ production at muon collider. The later process is considered both in the context of SM and in presence of the portal VLL extension of the SM.
The $\mu^{+}\mu^{-}\to h\gamma_{d}$ process is highly sensitive to the following independent model parameters: the dark photon mass ($m_{\gamma_{d}}$), dark gauge coupling ($g_{d}$), and both fermion ($\sin\theta_{L}$) and scalar ($\sin\theta_{s}$) mixing angles. We explore the non-decoupling effect owing to the presence of the muon-heavy muon-dark photon vertex and show that a significant region of parameter space consistent with the perturbative unitarity, electroweak precision measurements, lepton flavor universality, allowed ranges of scalar mixing angle and other relevant constraints can be found where the $h\gamma_{d}$ cross-section is $\sim$ few fb to $10^{2}$ fb. Consequently, one can achieve up to $10\%$ better precision in the measurement of Higgs properties w.r.t. that achievable in the $hZ$ channel.
A crucial outcome of our study is that this process not only provides an enhanced Higgs production rate at the muon collider but also helps to set a strong limit on the dark photon mass even for a fermion mixing angle as low as $10^{-6}$, independent of the gauge kinetic mixing parameter.
We have proposed two final states, i.e., leptonically quiet exactly one and exactly two central-jet(s) along with missing energy. These final states originate from the $h\gamma_{d}$ process when the Higgs boson decays to a pair of bottom quarks and the dark photon is treated as an invisible particle at the detector. In order to set a $2\sigma$ exclusion limit in the $m_{\gamma_{d}}-m_{\mu_{p}}$ plane using the $h\gamma_{d}$ process in the aforementioned final states, we have also considered an exhaustive study of the SM background processes which contribute to these final states. With the help of certain kinematic variables such as missing energy, requirement of the presence of substructure in the leading central CA1.2 jet, and the invariant mass associated with the $b\bar{b}$ system to lie around the Higgs mass it is possible to substantially reduce the SM background contributions and efficiently isolate the signal of $h\gamma_{d}$ production at the future muon collider.
[…]
Acknowledgement
We gratefully acknowledge Satyanarayan Mukhopadhyay for his valuable insights and useful comments. KT thanks the Department of Science and Technology (DST), Government of India, for financial support under the INSPIRE Fellowship Programme (IF220170). KT is also grateful to Sujay Mondal for providing access to high-performance computational facilities which significantly aided the progress of this work. KT further extends her thanks to the organizers and instructors of the Iwate Collider School 2025 for their excellent lectures and tutorial sessions. SV acknowledges the Physical Research Laboratory, Ahmedabad, for supporting him through the Post-Doctoral Fellowship.
Appendix A Vertex factors
In this appendix, we present Feynman rules for the relevant vertices assuming all the associated four momenta are coming out of the vertex. Below the $L$ and $R$ represent the left- and right- chiral fermions, respectively.
A.1 Vertices with Dark Photon:
  1. 1.
    $\bm{\mu^{+}_{p_{{}_{R}}}-\mu^{-}_{L}-\gamma_{d}}\penalty 10000\ /\penalty 10000\ \bm{\mu^{+}_{R}-\mu^{-}_{p_{{}_{L}}}-\gamma_{d}}$
    $\text{VF}_{1}$ $=$ $-ig_{d}\sin\theta_{L}\cos\theta_{L}\gamma^{\mu}$ (A.1)
  2. 2.
    $\bm{\mu^{+}_{p_{{}_{L}}}-\mu^{-}_{R}-\gamma_{d}}\penalty 10000\ /\penalty 10000\ \bm{\mu^{+}_{L}-\mu^{-}_{p_{{}_{R}}}-\gamma_{d}}$
    $\text{VF}_{2}$ $=$ $-ig_{d}\sin\theta_{R}\cos\theta_{R}\gamma^{\mu}$ (A.2)
  3. 3.
    $\bm{\mu^{+}_{{L}}-\mu^{-}_{R}-\gamma_{d}}$
    $\text{VF}_{3}$ $=$ $-ig_{d}\sin^{2}\theta_{R}\gamma^{\mu}$ (A.3)
  4. 4.
    $\bm{\mu^{+}_{{R}}-\mu^{-}_{L}-\gamma_{d}}$
    $\text{VF}_{4}$ $=$ $-ig_{d}\sin^{2}\theta_{L}\gamma^{\mu}$ (A.4)
A.2 Vertices with $Z$ boson:
  1. 1.
    $\bm{\mu^{+}_{p_{{}_{R}}}-\mu^{-}_{L}-Z}\penalty 10000\ /\penalty 10000\ \bm{\mu^{+}_{R}-\mu^{-}_{p_{{}_{L}}}-Z}$
    $\text{VF}_{5}$ $=$ $-\dfrac{i}{2}g_{z}\sin\theta_{L}\cos\theta_{L}\gamma^{\mu}$ (A.5)
  2. 2.
    $\bm{\mu^{+}_{{L}}-\mu^{-}_{R}-Z}$
    $\text{VF}_{6}$ $=$ $ig_{z}\sin^{2}\theta_{W}\gamma^{\mu}$ (A.6)
  3. 3.
    $\bm{\mu^{+}_{{R}}-\mu^{-}_{L}-Z}$
    $\text{VF}_{7}$ $=$ $i\penalty 10000\ \dfrac{g_{z}}{2}(2\sin^{2}\theta_{W}-\cos^{2}\theta_{L})\gamma^{\mu}$ (A.7)
A.3 Vertices with Higgs:
  1. 1.
    $\bm{\mu^{+}_{p_{{}_{L}}}-\mu^{-}_{L}-h}\penalty 10000\ /\penalty 10000\ \bm{\mu^{+}_{R}-\mu^{-}_{p_{{}_{R}}}-h}$
    $\text{VF}_{8}$ $=$ $-\dfrac{i}{\sqrt{2}}\sin\theta_{R}(\sin\theta_{L}\sin\theta_{s}\omega_{f}+\cos\theta_{L}\cos\theta_{s}y_{m})$ (A.8)
  2. 2.
    $\bm{\mu^{+}_{p_{{}_{R}}}-\mu^{-}_{R}-h}\penalty 10000\ /\penalty 10000\ \bm{\mu^{+}_{L}-\mu^{-}_{p_{{}_{L}}}-h}$
    $\text{VF}_{9}$ $=$ $\dfrac{i}{\sqrt{2}}\cos\theta_{R}(\cos\theta_{L}\sin\theta_{s}\omega_{f}-\sin\theta_{L}\cos\theta_{s}y_{m})$ (A.9)
  3. 3.
    $\bm{\gamma_{d}-\gamma_{d}-h}$
    $\text{VF}_{10}$ $=$ $-2ig_{d}m_{\gamma_{d}}\sin\theta_{s}$ (A.10)
  4. 4.
    $\bm{Z-Z-h}$
    $\text{VF}_{11}$ $=$ $ig_{z}m_{z}\cos\theta_{s}$ (A.11)
Appendix B Squared matrix element for $\mu^{+}\mu^{-}\to h\,\gamma_{d}$
The $t$- and $u$- channel diagrams for the $\mu^{+}\mu^{-}\to h\,\gamma_{d}$ process via portal vector-like lepton mediation ($\mu_{p}$) are presented in Fig 20.
Figure 20: Feynman diagrams for the process $\mu^{+}\mu^{-}\to h\,\gamma_{d}$ mediated by the portal matter $m_{\mu_{p}}$ through the (a) $t$-channel and (b) $u$-channel.
The amplitude for the $t$-channel diagram is given by
[…]
where $c_{s}^{t},\penalty 10000\ c_{p}^{t},\penalty 10000\ c_{{}_{V}}^{t}$ and $c_{{}_{A}}^{t}$ are given by
$c_{s}^{t}$ $=$ $\dfrac{m_{\mu_{p}}}{4}\sin{2\theta_{L}}\big[\dfrac{1}{D^{2}}\dfrac{g_{d}}{m_{\gamma_{d}}}\sin{\theta_{s}}(1-\dfrac{m^{2}_{\mu}}{m^{2}_{\mu_{p}}})(\sin^{2}{\theta_{L}}-\dfrac{m_{\mu}}{m_{\mu_{p}}}\cos^{2}{\theta_{L}})\penalty 10000\ +\penalty 10000\ \dfrac{1}{v_{{}_{\rm{EW}}}}\cos{\theta_{s}}(1+\dfrac{m_{\mu}}{m_{\mu_{p}}})\big]$
$c_{p}^{t}$ $=$ $-\dfrac{m_{\mu_{p}}}{4}\sin{2\theta_{L}}\big[\dfrac{1}{D^{2}}\dfrac{g_{d}}{m_{\gamma_{d}}}\sin{\theta_{s}}(1-\dfrac{m^{2}_{\mu}}{m^{2}_{\mu_{p}}})(\sin^{2}{\theta_{L}}+\dfrac{m_{\mu}}{m_{\mu_{p}}}\cos^{2}{\theta_{L}})\penalty 10000\ +\penalty 10000\ \dfrac{1}{v_{{}_{\rm{EW}}}}\cos{\theta_{s}}(1-\dfrac{m_{\mu}}{m_{\mu_{p}}})\big]$
$c_{{}_{V}}^{t}$ $=$ $\dfrac{g_{d}}{2}\cos{\theta_{L}}\sin{\theta_{L}}\big(\dfrac{m_{\mu}}{m_{\mu_{p}}}\dfrac{1}{D^{2}}+1\big)$
$c_{{}_{A}}^{t}$ $=$ $\dfrac{g_{d}}{2}\cos{\theta_{L}}\sin{\theta_{L}}\big(\dfrac{m_{\mu}}{m_{\mu_{p}}}\dfrac{1}{D^{2}}-1\big)$
The amplitude of the $u$-channel diagram is given by
[…]
where $c_{s}^{u},\penalty 10000\ c_{p}^{u},\penalty 10000\ c_{{}_{V}}^{u}$ and $c_{{}_{A}}^{u}$ are given by
$c_{s}^{u}$ $=$ $\dfrac{m_{\mu_{p}}}{4}\sin{2\theta_{L}}\big[\dfrac{1}{D^{2}}\dfrac{g_{d}}{m_{\gamma_{d}}}\sin{\theta_{s}}(1-\dfrac{m^{2}_{\mu}}{m^{2}_{\mu_{p}}})(\sin^{2}{\theta_{L}}-\dfrac{m_{\mu}}{m_{\mu_{P}}}\cos^{2}{\theta_{L}})\penalty 10000\ +\penalty 10000\ \dfrac{1}{v_{{}_{\rm{EW}}}}\cos{\theta_{s}}(1+\dfrac{m_{\mu}}{m_{\mu_{p}}})\big]$
$c_{p}^{u}$ $=$ $\dfrac{m_{\mu_{p}}}{4}\sin{2\theta_{L}}\big[\dfrac{1}{D^{2}}\dfrac{g_{d}}{m_{\gamma_{d}}}\sin{\theta_{s}}(1-\dfrac{m^{2}_{\mu}}{m^{2}_{\mu_{p}}})(\sin^{2}{\theta_{L}}+\dfrac{m_{\mu}}{m_{\mu_{P}}}\cos^{2}{\theta_{L}})\penalty 10000\ +\penalty 10000\ \dfrac{1}{v_{{}_{\rm{EW}}}}\cos{\theta_{s}}(1-\dfrac{m_{\mu}}{m_{\mu_{p}}})\big]$
$c_{{}_{V}}^{u}$ $=$ $\dfrac{g_{d}}{2}\cos{\theta_{L}}\sin{\theta_{L}}\big(\dfrac{m_{\mu}}{m_{\mu_{P}}}\dfrac{1}{D^{2}}+1\big)$
$c_{{}_{A}}^{u}$ $=$ $\dfrac{g_{d}}{2}\cos{\theta_{L}}\sin{\theta_{L}}\big(\dfrac{m_{\mu}}{m_{\mu_{P}}}\dfrac{1}{D^{2}}-1\big)$
The spin-averaged squared amplitude, neglecting the s channel dark photon mediated and t/u channel muon mediated contribution, is
$\overline{|\mathcal{M}|^{2}}$ $=$ $\overline{|\mathcal{M}_{t}|^{2}}+\overline{|\mathcal{M}_{u}|^{2}}+2\,\operatorname{Re}\!\left(\overline{\mathcal{M}_{t}\mathcal{M}_{u}^{*}}\right)$ (B.3)
where the $t$-channel contribution in the limit $m_{\mu}\to 0$ is
$\overline{|\mathcal{M}_{t}|^{2}}$ $=$ $\frac{2}{\left[\left(k_{1}-p_{1}\right)^{2}-m^{2}_{\mu_{p}}\right]^{2}}\Bigg[A_{1}\left\{2\left(k_{1}\cdot p_{1}\right)\left(k_{2}\cdot p_{1}\right)-p_{1}^{2}\left(k_{1}\cdot k_{2}\right)\right\}+A_{2}\penalty 10000\ m_{\mu_{p}}^{2}\left(k_{1}\cdot k_{2}\right)$ (B.4)
$+\frac{1}{m_{\gamma_{d}}^{2}}\Bigg(A_{1}\Big\{2\left(k_{1}\cdot p_{1}\right)\left(k_{2}\cdot p_{2}\right)\left(p_{1}\cdot p_{2}\right)-p_{1}^{2}\left(k_{1}\cdot p_{2}\right)\left(k_{2}\cdot p_{2}\right)-p_{2}^{2}\left(k_{1}\cdot p_{1}\right)\left(k_{2}\cdot p_{1}\right)$
$+\frac{p_{1}^{2}\penalty 10000\ p_{2}^{2}}{2}\left(k_{1}\cdot k_{2}\right)\Big\}+A_{2}\penalty 10000\ m_{\mu_{p}}^{2}\left\{\left(k_{1}\cdot p_{2}\right)\left(k_{2}\cdot p_{2}\right)-\frac{p_{2}^{2}}{2}\left(k_{1}\cdot k_{2}\right)\right\}\Bigg)\Bigg]$
Similarly, the $u$- channel contribution in the same limit reads
$\overline{|\mathcal{M}_{u}|^{2}}$ $=$ $\frac{2}{\left[\left(p_{1}-k_{2}\right)^{2}-m^{2}_{\mu_{p}}\right]^{2}}\Bigg[A_{1}\left\{2\left(k_{1}\cdot p_{1}\right)\left(k_{2}\cdot p_{1}\right)-p_{1}^{2}\left(k_{1}\cdot k_{2}\right)\right\}+A_{2}\penalty 10000\ m_{\mu_{p}}^{2}\left(k_{1}\cdot k_{2}\right)$ (B.5)
$+\frac{1}{m_{\gamma_{d}}^{2}}\Bigg(A_{1}\Big\{2\left(k_{1}\cdot p_{2}\right)\left(k_{2}\cdot p_{1}\right)\left(p_{1}\cdot p_{2}\right)-p_{1}^{2}\left(k_{1}\cdot p_{2}\right)\left(k_{2}\cdot p_{2}\right)-p_{2}^{2}\left(k_{1}\cdot p_{1}\right)\left(k_{2}\cdot p_{1}\right)$
$+\frac{p_{1}^{2}\penalty 10000\ p_{2}^{2}}{2}\left(k_{1}\cdot k_{2}\right)\Big\}+A_{2}\penalty 10000\ m_{\mu_{p}}^{2}\left\{\left(k_{1}\cdot p_{2}\right)\left(k_{2}\cdot p_{2}\right)-\frac{p_{2}^{2}}{2}\left(k_{1}\cdot k_{2}\right)\right\}\Bigg)\Bigg]$
Here, the coefficients $A_{1}$ and $A_{2}$ are as follows
$A_{1}$ $=$ $\left(c_{{}_{V}}^{2}+c_{{}_{A}}^{2}\right)\left(c_{s}^{2}+c_{p}^{2}\right)-4c_{s}\penalty 10000\ c_{p}\penalty 10000\ c_{{}_{V}}c_{{}_{A}}$ (B.6)
$A_{2}$ $=$ $\left(c_{{}_{V}}^{2}+c_{{}_{A}}^{2}\right)\left(c_{s}^{2}+c_{p}^{2}\right)+4c_{s}\penalty 10000\ c_{p}\penalty 10000\ c_{{}_{V}}c_{{}_{A}}$
with the coupling identifications
$c_{s}=c_{s}^{t}=c_{s}^{u},\qquad c_{p}=c_{p}^{t}=-\,c_{p}^{u},\qquad c_{V}=c_{V}^{t}=c_{V}^{u},\qquad c_{A}=c_{A}^{t}=c_{A}^{u}.$
The corresponding interference term is
$\overline{\mathcal{M}_{t}\mathcal{M}_{u}^{*}}$ $=$ $\frac{2}{\left[\left(k_{1}-p_{1}\right)^{2}-m_{\mu_{p}}^{2}\right]\left[\left(p_{1}-k_{2}\right)^{2}-m_{\mu_{p}}^{2}\right]}\Bigg[2A_{3}\left(k_{1}\cdot p_{1}\right)\left(k_{2}\cdot p_{1}\right)+A_{2}\penalty 10000\ m_{\mu_{p}}^{2}\left(k_{1}\cdot k_{2}\right)$ (B.7)
$-\frac{1}{m_{\gamma_{d}}^{2}}\Bigg(A_{3}\Big\{p_{1}^{2}\left(k_{1}\cdot p_{2}\right)\left(k_{2}\cdot p_{2}\right)+p_{2}^{2}\left(k_{1}\cdot p_{1}\right)\left(k_{2}\cdot p_{1}\right)-\left(p_{1}\cdot p_{2}\right)\left(k_{1}\cdot p_{2}\right)\left(k_{2}\cdot p_{1}\right)$
$-\left(p_{1}\cdot p_{2}\right)\left(k_{1}\cdot p_{1}\right)\left(k_{2}\cdot p_{2}\right)+\left(p_{1}\cdot p_{2}\right)^{2}\left(k_{1}\cdot k_{2}\right)-\frac{p_{1}^{2}\penalty 10000\ p_{2}^{2}}{2}\left(k_{1}\cdot k_{2}\right)\Big\}$
$+\frac{A_{2}}{2}m_{\mu_{p}}^{2}\left\{p_{2}^{2}\left(k_{1}\cdot k_{2}\right)-\left(k_{1}\cdot p_{2}\right)\left(k_{2}\cdot p_{2}\right)\right\}\Bigg)\Bigg]$
where the coefficient $A_{3}$ takes the form
$A_{3}$ $=$ $(c^{2}_{s}-c^{2}_{p})(c_{{}_{V}}^{2}-c_{{}_{A}}^{2})$ (B.8)
In the center-of-mass frame and in the limit $m_{h},\penalty 10000\ m_{\gamma_{d}}<<E_{{}_{\rm{CM}}}$, the $\overline{|\mathcal{M}|^{2}}$ for the pure $t$- and $u$- channels and the corresponding interference contribution take the form
$\overline{|\mathcal{M}_{t}|^{2}}$ $=$ $\frac{E_{{}_{\rm CM}}^{2}}{\left(a-b\cos\theta\right)^{2}}\Bigg[A_{1}\left(\frac{E_{{}_{\rm CM}}^{2}}{4}\sin^{2}\theta-m_{h}^{2}\right)+A_{2}\penalty 10000\ m_{\mu_{p}}^{2}+\frac{1}{2m_{\gamma_{d}}^{2}}\Bigg({A_{1}}\penalty 10000\ \Bigg\{\frac{E_{{}_{\rm CM}}^{4}}{4}\penalty 10000\ (1-\cos\theta)^{2}$ (B.9)
$-\frac{\left(m_{h}^{2}+m_{\gamma_{d}}^{2}\right)E_{{}_{\rm CM}}^{2}}{4}\sin^{2}\theta+m_{h}^{2}\penalty 10000\ m_{\gamma_{d}}^{2}\Bigg\}+A_{2}\penalty 10000\ {m_{\mu_{p}}^{2}}\left\{\frac{E_{{}_{\rm CM}}^{2}}{4}\sin^{2}\theta-m_{\gamma_{d}}^{2}\right\}\Bigg)\Bigg]$
$\overline{|\mathcal{M}_{u}|^{2}}$ $=$ $\frac{E_{{}_{\rm CM}}^{2}}{(a+b\cos\theta)^{2}}\Bigg[A_{1}\left(\frac{E_{{}_{\rm CM}}^{2}}{4}\sin^{2}\theta-m_{h}^{2}\right)+A_{2}\penalty 10000\ m_{\mu_{p}}^{2}+\frac{1}{2m_{\gamma_{d}}^{2}}\Bigg({A_{1}}\Bigg\{\frac{E_{{}_{\rm CM}}^{4}}{4}(1+\cos\theta)^{2}$ (B.10)
$-\frac{\left(m_{h}^{2}+m_{\gamma_{d}}^{2}\right)E_{{}_{\rm CM}}^{2}}{4}\sin^{2}\theta+m_{h}^{2}m_{\gamma_{d}}^{2}\Bigg\}+\,A_{2}\,{m_{\mu_{p}}^{2}}\left\{\frac{E_{{}_{\rm CM}}^{2}}{4}\sin^{2}\theta-m_{\gamma_{d}}^{2}\right\}\Bigg)\Bigg]$
$\overline{\mathcal{M}_{t}\mathcal{M}_{u}^{*}}$ $=$ $\frac{{E_{{}_{\rm CM}}^{2}}}{(a^{2}-b^{2}\cos^{2}\theta)}\Bigg[\left(A_{3}\frac{E_{{}_{\rm CM}}^{2}}{4}\sin^{2}\theta+A_{2}\penalty 10000\ m_{\mu_{p}}^{2}\right)+\frac{1}{2m_{\gamma_{d}}^{2}}\Bigg({A_{3}}\Bigg\{m^{2}_{h}m^{2}_{\gamma_{d}}$ (B.11)
$-\frac{\left(E_{{}_{\rm CM}}^{2}+m_{h}^{2}+m_{\gamma_{d}}^{2}\right){E_{{}_{\rm CM}}^{2}}}{4}\sin^{2}\theta\Bigg\}+A_{2}\penalty 10000\ m_{\mu_{p}}^{2}\left\{\frac{E_{{}_{\rm CM}}^{2}}{8}\sin^{2}\theta-m_{\gamma_{d}}^{2}\right\}\Bigg)\Bigg]$
Here,
$a=\frac{E_{\rm CM}^{2}}{2}+m_{\mu_{p}}^{2},\qquad b=\frac{E_{\rm CM}^{2}}{2}.$